An object's motion is described by the equation: d=5cos(pi/3(t))
The displacement d, is measured in meters. The time t, is measured in seconds.
Answer the following questions, showing all work:
(a) What is the object's position at t=0?
(b) What is the object's maximum displacement from its equilibrium position?
(c) How much time is required for one oscillation?
(d) At what time will the object first reach its equilibrium position (hint: d=0)?

Found 2 solutions by Alan3354, Boreal:
Answer by Alan3354(69443)

An object's motion is described by the equation: d=5cos(pi/3(t))
The displacement d, is measured in meters. The time t, is measured in seconds.
Answer the following questions, showing all work:
----
t is in the numerator.
d=5cos(pi*t/3)
----
(a) What is the object's position at t=0?
Sub zero for t.
-----
(b) What is the object's maximum displacement from its equilibrium position?
The amplitude, 5 meters.
-----
(c) How much time is required for one oscillation?
6 seconds.
P = 2pi/(pi/3) = 6
-------
(d) At what time will the object first reach its equilibrium position (hint: d=0)?
d=5cos(pi*t/3) = 0
cos(pi*t/3) = 0 = cos(pi/2)
pi*t/3 = pi/2
t/3 = 1/2
t = 1.5 seconds

Answer by Boreal(15235)

At t=0, the function is d=5 cos (0)=5
The maximum displacement is the cosine coefficient or 5.
For one oscillation, the period, it is 2pi/(pi/3)=6
For the equilibrium position, it is one-quarter of the period, or 1.5 pi/3
and 4.5 pi/3. The cosine is 1 at 0, 0 at pi/2 and again at 3pi/2